# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 3.1 Training Data Extraction
In Training Data Extraction attacks, an adversary aims to reconstruct training data having only access to a target Language Model. In particular, those attacks aim to extract the entire example by querying the target model (Carlini et al., 2021). Retrieval of examples is based on the assumption that they were memorized by the target models, and the definitions of memorization underlying these attacks may vary. However, the various definitions revolve around the key concept that a textual example or part of it can be extracted from a language model. In the original definition of Carlini et al. (2021), an example is defined to be extractable if there is a prefix used to condition the generation of the model so that the string can be exactly generated from the model. This definition is first introduced by Carlini et al. (2021) and then largely adopted with slight modifications (Huang et al., 2022; Carlini et al., 2023a; Nasr et al., 2023).
Training Data Extraction Extracting Training Data consists of conditioning a LM to the right prompt so that it is likely that it will generate a training example to complete it. Formally, a textual example $x$ is extractable from a LM $g$ if there exists a prompt $p$ such that: $x=\arg\max_{x^{\prime}}g(x^{\prime}|p)$. Once the right prompt has been identified, this procedure can be used to reconstruct training data, and hence, it can be used to extract sensitive information like Personally Identifiable Information (PII) from a model.
@@ 3.1.1 Non-adversarial extraction
Using training data extraction techniques, it has been shown that Language Models are prone to reveal private information when prompted with sentences that require the disclosure of such information. While trying to recover memorized training examples, Carlini et al. (2021) demonstrated that GPT-2 tends to memorize personal information such as Twitter handles, emails and UUIDs (Universal Unique Identifiers). In particular, in their experiments the language model generation is conditioned to some prompts that the attacker predefined: to maximize the probability of success of the attack, they choose prompts that are likely to be in the training data, like samples from the Common Crawl  [^2] https://commoncrawl.org/. Then, once the generation is complete, the attacker tries to figure out which sequences are most likely to be generated as a result of the memorization, that is, which of the generated strings exactly matches the training data. In order to identify sequences that are more likely to be generated because of memorization, Carlini et al. (2021) compare the likelihood that is assigned to each sequence by two models: one is the model under attack – GTP-2 – and the other is a reference model, another LM. The attacker retains as possibly memorized sentences only those that have a high likelihood in the target model when compared to the likelihood assigned by the reference model. This is a form of Membership Inference Attack §3.2. Finally, once the attacker has retained a subset of strings that are likely to be generated due to memorization, the success of the attack must be verified by checking whether these high-likelihood sequences are actually part of the training set. However, for this step, Carlini et al. (2021) could only manually verify the effectiveness of the attack by searching the Internet for matching sequences, as the tested model did not have accessible training data.
@@
The choice of the right prompt is crucial in this setting: in fact, when prompted with a prefix already seen during training, models tend to complete the prompt with the rest of the training sequence (Carlini et al., 2023a). Huang et al. (2022) further explore this direction and show that conditioning a model to a prompt already observed during training can lead to a leakage of PII, like emails, and that is more effective than designing new, previously unseen prompts. In particular, they distinguish between memorization and association. A model memorizes personal information if there exists a prompt – already part of the training data for that model – that leads the model to generate the personal information. The association requires that there exists a prompt – not necessarily observed during training and containing some reference to an individual – that can lead to the generation of PII by the model. Huang et al. (2022) show that the former approach is more effective than the latter, quantifying the number of e-mails that a model from the GPT-Neo family can predict when conditioned by the two different kinds of prompts.
@@
Those attacks can also be used to reconstruct the dataset used to fine-tune the model. To reconstruct personal information from a generative language model like GPT-2, Lukas et al. (2023) propose to solve a masked language modeling task to reconstruct masked personal information from a sentence, conditioning a model on both the prefix and suffix of the masked information: they call this approach PII Reconstruction. However, it requires the attacker to gain access to masked training data before performing the attack. Panchendrarajan & Bhoi (2021) reconstruct the training data of a fine-tuned GPT-2. This work stems from the observation – shared also by Membership Inference Attacks as described in Section 3.2 – that sentences that have been used in the fine-tuning phase will be given different probabilities by the fine-tuned version of a model and by the base model. They generate candidate sentences by starting the generation of the fine-tuned model with an empty string. Then they observe the generation of the base model over the same sentence. Results show that it is possible to reconstruct about 75% of informative words in the fine-tuning dataset.
@@
These experiments were the foundation for exploring other datasets and larger models. Nasr et al. (2023) demonstrate that the method proposed by Carlini et al. (2021) is more effective than initially thought: querying open-source models, they were able to verify the success of the attack procedure by accessing the training data only to verify the attack. They performed the attacks on models belonging to GPT-Neo, Pythia and RedPajama-INCITE families and using as initial prompts samples from Wikipedia. Wang et al. (2024) explore the leakage of training data of OpenAI’s GPT closed models evaluating the accuracy of information extraction of sensitive information contained in pre-training data. The private information they are able to extract is emails from the Enron Email dataset (Klimt & Yang, 2004). They hypothesize that this dataset is used during the training phase of GPT-3.5 and GPT-4 since it is included in the Pile dataset (Gao et al., 2020): while no information is available on the training set of those models, the Pile was used in the pre-training of the GPT-Neo family of models and hence, possibly, also in the closed models under investigation. They adopt the same prompts proposed by Huang et al. (2022) and show results also in few-shots prompting settings. The experiments demonstrate that GPT-3.5 and GPT-4 can predict, respectively, around 5% and 4% of the email addresses accurately. Moreover, in the few-shot prompting setting with emails extracted from the same dataset, these models accurately predict up to 48.19% of emails correctly. The same trend is not observed for prompts that contain emails that have a different domain from the target one, suggesting the importance of the right choice of demonstrations for training data extraction attacks. While LLMs are generally more resistant to attacks where association is required, the work of Shao et al. (2023) shows that it is possible to recover 3% of the training emails from a 20 billion parameter model with attacks that require association, meaning that the extraction procedure does not rely on prompts already observed during training. It has also been proposed to extract targeted examples from a model pre-training data: Zhang et al. (2023b) propose an attack to recover a certain suffix given a precise prefix, known to be in the pre-training data. They suppose that the attacker can also obtain some pairs of prefixes and suffixes and train an attack model on them: they tune soft prompt embeddings and then estimate the confidence that a generated suffix is, in fact, as it is in the corresponding prompt in the pre-training data. A similar approach, based on soft prompt tuning, is also adopted by Ozdayi et al. (2023). As the size of the models increases, this type of attack appears to be more effective (Nasr et al., 2023). Moreover, the same techniques have been also discussed for other models, like diffusion models (Carlini et al., 2023b).
@@
As concerns about the potential leakage of private information via training data extraction attacks grow, a number of tools are being investigated to assess and potentially prevent the leakage of private information by LMMs before it spreads. On the evaluation side, ProPILE (Kim et al., 2023) is a tool designed to test models with black-box and white-box attacks against their tendency to release PII: the black-box approach is designed to allow data subjects to test whether their PII is leaked; the white-box approach is based on soft prompting and allows providers to conduct their model surveys to prevent leaks of private information. Differential privacy (see §5 for further details) is also effective against the memorization of private information and can successfully defend against these attacks. Downey et al. (2022) demonstrate that by hand-crafting canary sentences containing sensitive content in the training data: the last two words of each canary are designed to contain the most sensitive information and the LM is asked to complete the canary. The GPT model they train with Differential Privacy – with some loss in model performance – successfully preserves privacy.
@@ 3.1.2 Adversarial prompting
There have been initiatives to ensure LLMs are aligned with human values and used appropriately. For example, the OpenAI usage policy requires that users do not use the models to compromise the privacy of others, including collecting or generating personal data. Moreover, OpenAI’s ChatGPT has been aligned with Reinforcement Learning from Human Feedback (Ouyang et al., 2022) and avoids answering questions regarding its training data. In this scenario, the techniques based on non-adversarial prompts may not be effective: Nasr et al. (2023) discuss how ChatGPT tends to not expose training data when prompted with prefixes that are likely seen during training. For these reasons, some attacks against LLMs define adversarial prompts that force the model to diverge from its originally intended use. For example, asking the model to repeat indefinitely a word, Nasr et al. (2023) notice that ChatGPT diverges and starts emitting training data, also containing personally identifiable information.
A new form of adversarial prompt attacks are jailbreak prompts. These prompts are designed to circumvent safeguards and manipulate LLMs into generating harmful content. In some cases, a goal hijacking is performed, meaning that the original goal of a prompt is changed to a new goal, like generating a target phrase, by injecting malicious text in the original prompt (Perez & Ribeiro, 2022). Shen et al. (2023) characterize a large set of jailbreak prompts and evaluate their effectiveness against different LLMs. They notice that most of these prompts encourage the model to act like another fictional assistant, such as DAN (short for Doing Anything Now), which has no limitation in the answers. Along with other potentially harmful scenarios, they show that this type of prompt allows a user to receive suggestions on how to collect personal information.
In this context, Li et al. (2023b) focused on the usage of adversarial prompts to pose privacy threats: they propose a multi-step jailbreak prompt to extract personal information, such as emails, based on Chain-of-Thought (Kojima et al., 2023). They study the effect of adversarial prompting against ChatGPT. They first notice that the model tends to avoid sharing private information if direct data extraction prompts are used. Then they propose the Multi-step Jailbreaking Prompt in three steps: they assume the role of the user to input the jailbreaking prompt, then act as the assistant to confirm that the jailbreak mode is enabled and, lastly, they play the role of the user and query the assistant with a non-adversarial prompt. They show that more frequent emails are less difficult to recover: experiments on the Enron dataset show that more than 50% of frequent emails in the dataset can be extracted.
@@ 3.2 Membership Inference Attacks (MIA)
Membership inference attacks (MIA) (Shokri et al., 2016) are currently a de facto standard when it comes to verifying model privacy (Murakonda & Shokri, 2020). In this kind of attack, the attacker has access to (1) some records (supposed to come from a similar distribution of the training data) and (2) to the model, which could be provided as an API, in a black box setting, or be fully available as a white box. The aim of the attacker is to find out if these records (1) are actually part of the training set for that model (2). The initial formulation of this attack by Shokri et al. (2016) required the attacker to build several "shadow models" that mimic the original model, trained on datasets that may resemble the distribution of the original one. Then, the attacker trains a classifier on the shadow models’ outputs to assess the usage of some records during the training phase by the shadow models (where the ground truth is known since the attacker also designed the training set of those models). The classifier, trained on shadow models’ outputs, is then used against the original model with the objective of identifying data that are part of its training set. The proposed method achieved an accuracy score of 90% on models trained on Google’s ML platform even without any prior assumption on the training set, showing that the attack is very effective. Ye et al. (2021) improve and expand upon this work by proposing several new MIAs and a framework to develop them. MIAs have also been shown to be effective in detecting whether a certain image was used in the training set of complex models for computer vision (Carlini et al., 2022).
@@
Differential Privacy is the best defence to implement against Membership Inference Attacks (see Section 5 for further details). Jagannatha et al. (2021) conduct various attacks against BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019) models trained on clinical data and show that Differential Privacy helps mitigate empirical privacy leakage. However, it has been shown that, if the privacy budget is too high, MIAs can also be effective against models trained with Differential Privacy (Truex et al., 2019; Nguyen et al., 2023).
3.2.1 MIA with Shadow Models
Stemming from the general setting described in Shokri et al. (2016), MIAs have been successfully tailored for Language Models to extract sensitive information from them. Song & Shmatikov (2019) argue that Language Models tend to memorize training data and, therefore, a user can tell whether their data have been used also in a black box setting. Their attack positively exploits the tendency of these models to rank relatively rare words higher when they are in the same context in which they were seen during training. Hisamoto et al. (2020) try to make this type of attack more general by training an attack classifier not on the output probabilities but only on features that can be extracted from the output sequences of a sequence-to-sequence Transformer model. Carlini et al. (2022) observe that MIAs could be better framed as the procedure of comparing models that are trained on a given example and models that are not trained on it, and to find a difference in their behaviour, for example having a likelihood of the example much higher in the first case rather than in the other. They train numerous shadow GPT-2 models to measure the probability of observing a certain likelihood of an example in models trained and not trained on it. However, training shadow models may be too expensive in the context of LLMs: the original formulation of MIA can be used as long as it is feasible for the attacker to construct shadow models that are similar to the target model.
@@ 3.2.2 MIA with Thresholds
Since the construction of shadow models can be too expensive, MIAs have also been formulated without relying on them. In fact, some MIAs have been designed to exploit some measure of overfit over the training data in order to understand whether a given example has already been seen by a model: overfitting the training data is a sufficient condition to detect memorized examples (Yeom et al., 2018). Shi et al. (2023) observe that the probability distribution is more spread over low probability words for an unseen example: they design attacks that threshold the average log-likelihood of the top-k rarest words to ascertain whether or not the example is part of the training data. Language models trained on clinical data have been shown to be sensitive to MIAs. Mireshghallah et al. (2022) demonstrate the effectiveness of MIAs designed against Language Models trained with Masked Language Modeling objectives, obtaining a substantial gain both in AUC and in recall of the attack over the baseline. Their method is based on thresholding the ratio between the likelihood of a sentence on the target model and the likelihood of the same sentence on a model trained on a general population. Mattern et al. (2023) argue that the limited availability of training data for the reference model poses a strong restriction on the applicability of previous work. They propose to generate perturbations of a sentence and to measure the average loss of a target model on the original sentence and on the perturbations: they hypothesize that, if the sentence is not in the training set, the model should behave similarly also on the perturbed examples, while it should have a lower loss conditioned on the original sentence if it is part of the training set. These kinds of attacks are also useful to understand, given two snapshots of a model – for example a pre-trained model and a fine-tuned version of it – on which data the fine-tuning phase was performed. In particular, the difference between the probability assigned from different snapshots to each token in a sentence can help identify whether an example was used during later training phases (Zanella-Béguelin et al., 2020). Duan et al. (2023b) show that LLMs, while being extremely versatile in a number of tasks thanks to their in-context capabilities, can be subject to MIAs that target datasets used during prompt training. They demonstrate that an LLM trained to answer to prompts tends to expose more information than the corresponding fine-tuned model.
3.3 Model Inversion

### v2

@@ 3.1 Training Data Extraction
In Training Data Extraction attacks, an adversary aims to reconstruct training data having only access to a target Language Model. In particular, those attacks aim to extract the entire example by querying the target model (Carlini et al., 2021). Retrieval of examples is based on the assumption that the target models memorized them, and the definitions of memorization underlying these attacks may vary. However, the various definitions revolve around the key concept that a textual example or part of it can be extracted from a language model. In the original definition of Carlini et al. (2021), an example is defined to be extractable if there is a prefix used to condition the generation of the model so that the string can be exactly generated from the model. This definition is first introduced by Carlini et al. (2021) and then largely adopted with slight modifications (Huang et al., 2022; Carlini et al., 2023a; Nasr et al., 2023).
Training Data Extraction Extracting Training Data consists of conditioning an LM to the right prompt so that it will likely generate a training example to complete it. Formally, a textual example $x$ is extractable from a LM $g$ if there exists a prompt $p$ such that: $x=\arg\max_{x^{\prime}}g(x^{\prime}|p)$. Once the right prompt has been identified, this procedure can be used to reconstruct training data, and hence, it can be used to extract sensitive information like Personally Identifiable Information (PII) from a model.
@@ 3.1.1 Non-adversarial extraction
Using training data extraction techniques, it has been shown that Language Models are prone to reveal private information when prompted with sentences requiring disclosing such information. While trying to recover memorized training examples, Carlini et al. (2021) demonstrated that GPT-2 tends to memorize personal information such as Twitter handles, emails, and UUIDs (Universal Unique Identifiers). In particular, in their experiments, the language model generation is conditioned to some prompts that the attacker predefined: to maximize the probability of success of the attack, they choose prompts that are likely to be in the training data, like samples from the Common Crawl  [^2] https://commoncrawl.org/. Then, once the generation is complete, the attacker tries to figure out which sequences are most likely to be generated as a result of the memorization, that is, which of the generated strings exactly matches the training data. In order to identify sequences that are more likely to be generated because of memorization, Carlini et al. (2021) compare the likelihood that is assigned to each sequence by two models: one is the model under attack – GTP-2 – and the other is a reference model, another LM. The attacker retains as possibly memorized sentences only those that have a high likelihood in the target model when compared to the likelihood assigned by the reference model. This is a form of Membership Inference Attack §3.2. Finally, once the attacker has retained a subset of strings that are likely to be generated due to memorization, the success of the attack must be verified by checking whether these high-likelihood sequences are actually part of the training set. However, for this step, Carlini et al. (2021) could only manually verify the effectiveness of the attack by searching the Internet for matching sequences, as the tested model did not have accessible training data.
@@
The choice of the right prompt is crucial in this setting: in fact, when prompted with a prefix already seen during training, models tend to complete the prompt with the rest of the training sequence (Carlini et al., 2023a). Huang et al. (2022) further explore this direction and show that conditioning a model to a prompt already observed during training can lead to a leakage of PII, like emails, and that is more effective than designing new, previously unseen prompts. In particular, they distinguish between memorization and association. A model memorizes personal information if a prompt exists- already part of the training data for that model- that leads the model to generate personal information. The association requires that a prompt – not necessarily observed during training and containing some reference to an individual – exists that can lead to the generation of PII by the model. Huang et al. (2022) show that the former approach is more effective than the latter, quantifying the number of e-mails that a model from the GPT-Neo family can predict when conditioned by the two different kinds of prompts. Moreover, they notice how these attacks get more effective as the size of the models increases, and – as also noted in as also noted in Carlini et al. (2023a) and Nasr et al. (2023)–larger models tend to be more exposed to training data extraction attacks.
@@
Those attacks can also be used to reconstruct the dataset used to fine-tune the model. To reconstruct personal information from a generative language model like GPT-2, Lukas et al. (2023) propose to solve a masked language modelling task to reconstruct masked personal information from a sentence, conditioning a model on both the prefix and suffix of the masked information: they call this approach PII Reconstruction. However, the attacker must gain access to masked training data before performing the attack. Panchendrarajan & Bhoi (2021) reconstruct the training data of a fine-tuned GPT-2. This work stems from the observation – also shared by Membership Inference Attacks as described in §3.2 – that sentences that have been used in the fine-tuning phase will be given different probabilities by the fine-tuned version of a model and by the base model. They generate candidate sentences by starting the generation of the fine-tuned model with an empty string. Then, they observe the generation of the base model over the same sentence. Results show that it is possible to reconstruct about 75% of informative words in the fine-tuning dataset.
@@
These experiments were the foundation for exploring other datasets and larger models. Nasr et al. (2023) demonstrate that the method proposed by Carlini et al. (2021) is more effective than initially thought: querying open-source models, they were able to verify the success of the attack procedure by accessing the training data only to verify the attack. They performed the attacks on models belonging to GPT-Neo, Pythia and RedPajama-INCITE families and using as initial prompts samples from Wikipedia. Wang et al. (2024a) explore the leakage of training data of OpenAI’s GPT closed models evaluating the accuracy of information extraction of sensitive information contained in pre-training data. The private information they are able to extract is emails from the Enron Email dataset (Klimt & Yang, 2004). They hypothesize that this dataset is used during the training phase of GPT-3.5 and GPT-4 since it is included in the Pile dataset (Gao et al., 2020): while no information is available on the training set of those models, the Pile was used in the pre-training of the GPT-Neo family of models and hence, possibly, also in the closed models under investigation. They adopt the same prompts proposed by Huang et al. (2022) and show results also in few-shots prompting settings. The experiments demonstrate that GPT-3.5 and GPT-4 can predict, respectively, around 5% and 4% of the email addresses accurately. Moreover, in the few-shot prompting setting with emails extracted from the same dataset, these models accurately predict up to 48.19% of emails correctly. The same trend is not observed for prompts that contain emails that have a different domain from the target one, suggesting the importance of the right choice of demonstrations for training data extraction attacks. While LLMs are generally more resistant to attacks where association is required, the work of Shao et al. (2024) shows that it is possible to recover 3% of the training emails from a 20 billion parameter model with attacks that require association, meaning that the extraction procedure does not rely on prompts already observed during training. It has also been proposed to extract targeted examples from a model pre-training data: Zhang et al. (2023b) propose an attack to recover a certain suffix given a precise prefix known to be in the pre-training data. They suppose that the attacker can also obtain some pairs of prefixes and suffixes and train an attack model on them. They tune soft prompt embeddings and then estimate the confidence that a generated suffix is, in fact, as it is in the corresponding prompt in the pre-training data. A similar approach, based on soft prompt tuning, is also adopted by Ozdayi et al. (2023). As the size of the models increases, this type of attack appears to be more effective (Nasr et al., 2023). Moreover, the same techniques have also been discussed for other models, like diffusion models (Carlini et al., 2023b).
@@
As concerns about the potential leakage of private information via training data extraction attacks grow, a number of tools are being investigated to assess and potentially prevent the leakage of private information by LMMs before it spreads. On the evaluation side, ProPILE (Kim et al., 2023) is a tool designed to test models with black-box and white-box attacks against their tendency to release PII: the black-box approach is designed to allow data subjects to test whether their PII is leaked; the white-box approach is based on soft prompting and allows providers to conduct their model surveys to prevent leaks of private information. Differential privacy (see §5 for further details) is also effective against the memorization of private information and can successfully defend against these attacks. Downey et al. (2022) demonstrate that by hand-crafting canary sentences containing sensitive content in the training data, the last two words of each canary are designed to contain the most sensitive information, and the LM is asked to complete the canary. The GPT model they train with Differential Privacy – with some loss in model performance – successfully preserves privacy.
@@ 3.1.2 Adversarial prompting
There have been initiatives to ensure LLMs are aligned with human values and used appropriately. For example, the OpenAI usage policy requires that users do not use the models to compromise the privacy of others, including collecting or generating personal data. Moreover, OpenAI’s ChatGPT has been aligned with Reinforcement Learning from Human Feedback (Ouyang et al., 2022) and avoids answering questions regarding its training data. In this scenario, the techniques based on non-adversarial prompts may not be effective: Nasr et al. (2023) discuss how ChatGPT tends not to expose training data when prompted with prefixes that are likely seen during training. For these reasons, some attacks against LLMs define adversarial prompts that force the model to diverge from its originally intended use. For example, asking the model to repeat a word indefinitely, Nasr et al. (2023) notice that ChatGPT diverges and starts emitting training data, also containing personally identifiable information.
A new form of adversarial prompt attacks is jailbreak prompts. These prompts are designed to circumvent safeguards and manipulate LLMs into generating harmful content and are largely used agaist closed LLMs alligned via RLHF. In some cases, a goal hijacking is performed, meaning that the original goal of a prompt is changed to a new goal, like generating a target phrase, by injecting malicious text in the original prompt (Perez & Ribeiro, 2022). Shen et al. (2024) characterize a large set of jailbreak prompts and evaluate their effectiveness against different LLMs. They notice that most of these prompts encourage the model to act like another fictional assistant, such as DAN (short for Doing Anything Now), which has no limitation in the answers. Along with other potentially harmful scenarios, they show that this type of prompt allows a user to receive suggestions on how to collect personal information.
In this context, Li et al. (2023c) focused on the usage of adversarial prompts to pose privacy threats: they propose a multi-step jailbreak prompt to extract personal information, such as emails, based on Chain-of-Thought (Kojima et al., 2024). They study the effect of adversarial prompting against ChatGPT. They first notice that the model tends to avoid sharing private information if direct data extraction prompts are used. Then, they propose the Multi-step Jailbreaking Prompt in three steps: they assume the role of the user to input the jailbreaking prompt, then act as the assistant to confirm that the jailbreak mode is enabled Lastly, they play the role of the user and query the assistant with a non-adversarial prompt. They show that more frequent emails are less difficult to recover: experiments on the Enron dataset show that more than 50% of frequent emails in the dataset can be extracted.
@@ 3.2 Membership Inference Attacks (MIA)
Membership inference attacks (MIA) (Shokri et al., 2017) are currently a de facto standard when it comes to verifying model privacy (Murakonda & Shokri, 2020). In this kind of attack, the attacker has access to (1) some records (supposed to come from a similar distribution of the training data) and (2) to the model, which could be provided as an API, in a black box setting, or be fully available as a white box. The aim of the attacker is to find out if these records (1) are actually part of the training set for that model (2). The initial formulation of this attack by Shokri et al. (2017) required the attacker to build several "shadow models" that mimic the original model, trained on datasets that may resemble the distribution of the original one. Then, the attacker trains a classifier on the shadow models’ outputs to assess the usage of some records during the training phase by the shadow models (where the ground truth is known since the attacker also designed the training set of those models). The classifier, trained on shadow models’ outputs, is then used against the original model with the objective of identifying data that are part of its training set.
@@
Stemming from the general setting described in Shokri et al. (2017), MIAs have been successfully tailored for Language Models to extract sensitive information from them (Song & Shmatikov, 2019; Hisamoto et al., 2020). Carlini et al. (2022) observe that MIAs could be better framed as the procedure of comparing models that are trained on a given example and models that are not trained on it, and to find a difference in their behaviour, for example, having a likelihood of the example much higher in the first case rather than in the other. They train numerous shadow GPT-2 models to measure the probability of observing a certain likelihood of an example in models trained and not trained on it.
However, training shadow models may be too expensive in the context of LLMs: the original formulation of MIA can be used as long as it is feasible for the attacker to construct shadow models that are similar to the target model. For this reason, in the context of LLM the formulation for MIA has been adapted, generally eliminating shadow model developments.
@@ 3.2.1 MIA with Thresholds
Since the construction of shadow models can be too expensive, MIAs have also been formulated without relying on them. In fact, some MIAs against LLMs have been designed to exploit some measure of overfit over the training data in order to understand whether a given example has already been seen by a model: overfitting the training data is a sufficient condition to detect memorized examples (Yeom et al., 2018). Shi et al. (2024) observe that the probability distribution is more spread over low probability words for an unseen example: they design attacks that threshold the average log-likelihood of the top-k rarest words to ascertain whether or not the example is part of the training data. Language models trained on clinical data have been shown to be sensitive to MIAs. Mireshghallah et al. (2022) demonstrate the effectiveness of MIAs designed against Language Models trained with Masked Language Modeling objectives, obtaining a substantial gain both in AUC and in recall of the attack over the baseline. Their method is based on thresholding the ratio between the likelihood of a sentence on the target model and the likelihood of the same sentence on a model trained on a general population. While their experiments are conducted on Transformer-based networks trained on Masked Language Modeling objective, the method proposed can be, at least in principle, generalized also to the generative context. Mattern et al. (2023) argue that the limited availability of training data for the reference model poses a strong restriction on the applicability of previous work. They propose to generate perturbations of a sentence and to measure the average loss of a target model on the original sentence and on the perturbations: they hypothesize that if the sentence is not in the training set, the model should behave similarly also on the perturbed examples, while it should have a lower loss conditioned on the original sentence if it is part of the training set. These kinds of attacks are also useful to understand, given two snapshots of a model – for example, a pre-trained model and a fine-tuned version of it – on which data the fine-tuning phase was performed. In particular, the difference between the probability assigned from different snapshots to each token in a sentence can help identify whether an example was used during later training phases (Zanella-Béguelin et al., 2020). Duan et al. (2023) show that LLMs, while being extremely versatile in a number of tasks thanks to their in-context capabilities, can be subject to MIAs that target datasets used during prompt training. They demonstrate that an LLM trained to answer prompts exposes more information than the corresponding fine-tuned model.
However, a recent body of research poses some critical challenges to the use of MIAs to assess their reliability in demonstrating the inclusion of a sample in LLM training data. In some cases, the theoretical foundation of MIA is discussed. Zhang et al. (2024c) argue that, since it is no longer possible to use shadow models that do not include the data point in their training set, the statistical framework behind MIA is no longer applicable since other approaches are statistically unsound. Duan et al. (2024b) demonstrate empirically that a range of MIA is not reliable against LLMs. According to their findings, training for a single epoch- as is becoming standard practice for LLMs- and increasing amounts of training data may negatively affect MIA accuracy. They also notice that MIA become more difficult as the difference between data included in the training set and data not included becomes more subtle. However, as training corpora become larger – their statistics are based on the Pile (Gao et al., 2020) – this similarity increases: for example, non-members Github samples share on average 76.9 % of their 7-grams with the Pile. Hence, the combined effect of training techniques that diverge from the one defined for other models and the huge training sets may challenge MIA’s effectiveness against LLMs.
3.3 Model Inversion and Stealing